OpenStudy (kpop4life123):

⦁ Find the length of the missing side. Round your answer to the nearest tenth. @tanabugg

OpenStudy (kpop4life123):



OpenStudy (tanabugg):

OKau

OpenStudy (anonymous):

whats the missing side

OpenStudy (tanabugg):

10

OpenStudy (tanabugg):

9*9= 81 4*4 = 16 81+16 = 97 square root of 97 is 9.8488... Round up to get 10

OpenStudy (kpop4life123):

thanks alot!

OpenStudy (tanabugg):

Sure! Do you have anymore?

OpenStudy (kpop4life123):

yes :)

OpenStudy (tanabugg):

Okay. I'll be happy to help!

OpenStudy (kpop4life123):

⦁ The hypotenuse of a right triangle is 12, and one of the legs is 4. Find the length of the missing leg to the nearest hundredth.

OpenStudy (tanabugg):

13

OpenStudy (anonymous):

I'd disagree. The prompt requested that you round to the nearest *tenth*. The answer would have to be 9.8 instead of 10 as the tenth decimal value is obviously the first placeholder after the period.

OpenStudy (tanabugg):

12*12= 144 4*4= 16 144+ 16= 160 square root of 160= 12.6... round up to 13

OpenStudy (tanabugg):

Oh! Sorry I wasn't paying attention to that!

OpenStudy (kpop4life123):

:)